value of x in: [sq.root(x)/sq.root(1-x)] + sq.root(1-x)=5/2 ? How to solve this?

value of x in:
[sq.root(x)/sq.root(1-x)] + sq.root(1-x)=5/2 ?

How to solve this?

Grade:

1 Answers

Arun Kumar IIT Delhi
9 years ago
see 5/2=2+1/2
and 1-x>0 => x<1
x>0
now part with [] has to be integer and sqrt(1-x) will be < less than 1 and will contribute the fractional part
so sqrt(x-1)=1/2
x=3/4
but par under [] has to be <3 and >2
so solving that we get x belongs (4/5,9/10)
so no possible solution